OpenStudy (anonymous):

I need to: Find the x-intercept and y-intercept. y = 2x + 6

OpenStudy (anonymous):

The y=intercept is 6

OpenStudy (anonymous):

for x-intercept, let y=0. 0=2x+6 -6 -6 -6=2x divided by 2 on each side -3=x The x-intercept is -3

OpenStudy (anonymous):

okay it was the y part that confused me.

OpenStudy (anonymous):

For x-intercept, let y=0 for y=intercept, let x=0. y=2(0)+6 y=6

OpenStudy (anonymous):

A store had 235 MP3 players in the month of January. Every month, 30% of the MP3 players were sold and 50 new MP3 players were stocked in the store. Which recursive function best represents the number of MP3 players in the store f(n) after n months?